The textfit package for scaling up text to a desired size^{∗}

Sebastian Rahtz  Phil Taylor

1994/04/15

Contents

1 Introduction 1

1 Introduction

The user needs to be able to specify the width or height for text, and use the largest size necessary to fill that width or height. You can choose whether to have the size increased by simple points, or by using magsteps. Usage:

\scaletowidth{3in}{This is the way of the world}\scaletoheight{7mm}{This is the way of the world}

Details are controlled by two options: ‘magstep’ will make the system use magsteps, not points, and ‘noisy’ will produce useful messages.

Examples; note that the base font is taken from the current state when you enter the macros. so if you start off with a large font (eg cmr17), it will give a different result from that obtained by scaling up cmr5.

This is the way of the world

This is the way of the world

This is the way of the world

This is the way of the world

Cowabunga!

∗This file has version number 5, last revised 1994/04/15.

Gloucestershire Warwickshire Railway

This is the way of the world

This is the way of the world

This is the way of the world

This is the way of the world

Cowabunga!